OpenStudy (anonymous):

how do you find the surface area of a sphere with a Radius of 15cm using this formula SA = 4πr 2

OpenStudy (zzr0ck3r):

replace r with 15

OpenStudy (mathstudent55):

\(SA = 4 \pi \color{red}r^2\) \(\color{red}{r = 15~cm} \)

OpenStudy (anonymous):

plug that ^^^ into your calculator....

OpenStudy (anonymous):

but i need to show work

OpenStudy (zzr0ck3r):

\(SA=4\pi r^2 \implies SA = 4\pi15^2= 4*\pi*150=150*4*\pi=900\pi cm^2\)

OpenStudy (zzr0ck3r):

you see?

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

15^2 = 225

OpenStudy (mathstudent55):

\(SA = 4 \pi \color{red}r^2\) \(SA = 4 \pi \color{red}{(15~cm)}^2\) \(SA = 4 \pi \times 225~cm^2\) \(SA = 900\pi ~cm^2\)